6. What is the recurrence relation for :

   a) heap sort

b) selection sort

c) Shell sort

d) quick sort

e) merge sort

f) insertion sort
Answer #1

Heap Sort : T(n) = T(n − 2) + log2(n − 1) + log2(n)

Selection Sort : T(n) = T(n-1) + n

Shell Sort : T(n) = T(2k - 1)

Quick Sort : T(n) = 2T(n / 2) + n

Merge Sort : T(n) = 2(n/2) + O(n)

Insertion Sort : T(n) = T(n-1) + n